The solubility of lead(II) fluoride in water is 0.0175% (w/v) at 29°C. Calculate the Ksp of this salt at this temperature

  1. 1.46 x 10-12 M
  2. 1.46 x 10-9 M
  3. 3.64 x 10-10 M
  4. 1.31 x 10-9 M
  5. 1.31 x 10-12 M
  6. 3.28 x 10-10 M
1
Expert's answer
2021-06-07T03:23:37-0400

Solution

Solubility PbF2=0.0175w/v

Molar mass = 245.2g/mol

PbF2(s)↔️ Pb2+(aq)+2F-(aq)

ksp=[Pb2+] [F-]2


Molar solubility= "\frac{0.0175}{L}\times\frac{1mol}{245.2}= 7.137\times10^{-5}"


Dissociation equation shows for every mole of PbF2 that dissociates, 1mol of Pb2+ and 2 mol of F- are produced


Pb2+= 7.137"\times10^{-5}"

F-= "2\times7.137\times10^{-5}
=1.427\times10^{-4}"

Ksp=(7.137"\times 10^{-5})(1.427\times10^{-4})" 2


=1.4533"\times 10^-12"

=1.46"\times 10^{-12}M" M